OpenStudy (anonymous):

Limit as x approaches 0 (2x/cos(x))

OpenStudy (zarkon):

plug in zero

OpenStudy (zarkon):

\[\frac{2\cdot 0}{\cos(0)}\]

OpenStudy (zarkon):

what does that give you